OpenStudy (anonymous):

one number is 3 times another. Their sum is 8. Find the numbers

OpenStudy (anonymous):

2 and 6

OpenStudy (anonymous):

x and 3x x+3x=8 x=2 so another number is 6

OpenStudy (anonymous):

let smaller no is x bigger no = 3x x + 3x = 8 4x = 8 x = 8/4 x = 2 so smaller no is 2 bigger no is 3(2) i.e. 6